OpenStudy (anonymous):

Which equation represents the line that passes through (–6, 7) and (–3, 6)?

OpenStudy (anonymous):

y = –x + 9 y = –x + 5 y = –3x – 11y y = –3x + 25

OpenStudy (taramgrant0543664):

y=mx+b Slope= m= (y2-y1)/(x2-x1)

OpenStudy (taramgrant0543664):

y2=6 Y1=7 X1=-6 X2=-3

OpenStudy (anonymous):

so it would be c

OpenStudy (anonymous):

none of those answers are correct

OpenStudy (taramgrant0543664):

@satellite73 i was thinking the same thing!

OpenStudy (anonymous):

slope is \(-\frac{1}{3}\) and none of your potential answer have that slope

OpenStudy (taramgrant0543664):

Slope should be -1/3 using the slope equation

OpenStudy (anonymous):

so it would be c

OpenStudy (taramgrant0543664):

Using the slope of -1/3 substituting values in for x and y shouldn't the equation be y=-1/3x+5?

OpenStudy (anonymous):

thanks